40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14975     Entry time: Thu Oct 17 12:34:51 2019
Author: gautam 
Type: Update 
Category: General 
Subject: Daytime wishlist 

Some ideas that would help increase the locking duty-cycle in the short term. 

  1. Seismometer investigation - something is not quite right with the vertex seismometer. This is the one that is primarily used for feedforward, and can be really helpful.
  2. Drifting TTs - it is really annoying to have to re-set the input pointing into the interferometer every ~ hour. See Attachment #1.
  3. FSS - this isn't a scientific statement, but there were ~20-30 minute periods last night where the PC drive RMS was displaying sharp spikes repeating every 2-3 seconds, first with increasing and then decreasing height. This is a new feature to me in the long standing PC drive saga but it doesn't tell me exactly what is going on as I don't know in what frequency band the glitch is actually happening. See Attachment #2.
  4. ALS noise - while it is possible now to routinely transition the arm length control from the POX/POY to CARM/DARM basis, I see some sharp (<0.1 s) dives in the TRX/TRY levels when the arms are under ALS control. This wasn't present a week ago. Needs to be investigated - I defer this to the daytime tomorrow.
Attachment 1: DriftingTTs.png  88 kB  Uploaded Thu Oct 17 13:44:37 2019  | Hide | Hide all
DriftingTTs.png
Attachment 2: FSSweirdness.png  37 kB  Uploaded Thu Oct 17 13:44:53 2019  | Hide | Hide all
FSSweirdness.png
ELOG V3.1.3-